The total charge that an automobile battery can supply without being recharged is given in terms of ampere-hours. A typical 12-V battery has a rating of 60 ampere-hours (60 A for 1 h, 30 A for 2 h, and so on). Suppose that you forget to turn the headlights off in your parked automobile.

If each of the two headlights draws 2.1 A, how long will it be before your battery is "dead"?
Answer #1

total current = 2.1 + 2.1 = 4.2 A

time required = I/rating

= 4.2 / 60

= 0.07 h = 0.07*60 = 4.2 minutes

so the battery will die in 4.2 minutes